x - 2 square root of x = 0
Solve for all values of x that satisfies the equation.

Answer by stanbon(48554)

x-2(sqrt x)=0
2sqrt x = x
Square both sides to get:
4x=x^2
x^2-4x=0
x(x-2)=0
x=0 or x=2
Check answers in the original equation.
x=0 works
x=2 does not.
Only solution is x=0
Cheers,
Stan H.